OpenStudy (anonymous):

2/3+5/9 _________ 1/4+1/12 simplify the complex fraction

OpenStudy (anonymous):

get a common denominators...: 2/3 +5/9=6/9 +5/9 = ... 1/4 +1/12 = 3/12 +1/12 =... can you solve this ones?